# Using Large Language Models for Template Detection from Security Event Logs

### v1

@@ Abstract
Index Terms:
LLM-based template detection from security event logs, template detection from event logs, security event log analysis
I Introduction
Event log analysis is an important cyber security monitoring technique. For example, modern Security Operations Center (SOC) platforms like Splunk [1] and ElasticStack [2] rely on collecting and analyzing event logs from a large number of sources, whereas dedicated event correlation tools like SEC [3] have been designed for real-time monitoring of event logs. Although during recent years several structured event log formats have been proposed, a significant part of event log messages remain unstructured or semi-structured. As an example, the widely used BSD syslog protocol [4] defines the priority, timestamp, hostname, and syslog tag (program name with its process ID) fields, whereas the message text, which is the most informative part of the syslog message, remains a free-form string. Figure 1 depicts some example syslog messages without the priority, timestamp, and hostname fields (note that all syslog messages have been wrapped in Figure 1 due to their length, and they appear in one line in real-life event logs).
# sshd successful authentication messages and a template sshd[12992]: Accepted password for william from 192.168.3.1 port 60676 ssh2 sshd[17837]: Accepted publickey for felix from 192.168.17.8 port 54097 ssh2 sshd[<*>]: Accepted <*> for <*> from <*> port <*> ssh2 # sshd failed authentication messages and a template sshd[21378]: Failed none for alice from 192.168.3.250 port 60172 ssh2 sshd[6771]: Failed password for felix from 192.168.2.39 port 59020 ssh2 sshd[<*>]: Failed <*> for <*> from <*> port <*> ssh2 # Suricata IDS alert messages and a template suricata[1815]: [1:2101411:13] GPL SNMP public access udp [Classification: Attempted Information Leak] [Priority: 2] {UDP} 192.168.130.19:55919 -> 192.168.153.226:161 suricata[19225]: [1:2034455:2] ET EXPLOIT Possible Gitlab CE/EE Image Parser RCE Inbound (CVE-2021-22205) [Classification: Attempted Administrator Privilege Gain] [Priority: 1] {TCP} 192.168.1.7:3801 -> 192.168.81.29:80 suricata[<*>]: [<*>] <*> [Classification: <*>] [Priority: <*>] {<*>} <*> -> <*>
Figure 1: Example syslog messages and their templates.
@@ I Introduction
During the last two decades, a number of template detection algorithms have been proposed [5, 6, 7, 8, 9, 11, 12, 13, 14, 10, 15, 17, 16, 18, 21, 20, 22, 19, 23], and several comparative studies have been published for these algorithms [25, 24, 26]. Despite a significant amount of research in the domain of template mining, some research gaps have remained. Firstly, template detection approaches that have not employed LLMs (we call them traditional algorithms in the remainder of this paper) are unsupervised data mining-based approaches which can be tuned through hyperparameters. However, as pointed out in [21], selecting optimal values for hyperparameters is often not an easy task. Secondly, LLM-based template detection approaches are generally supervised and require labeled training data [19, 20, 22, 21]. As discussed in [23], the creation of such training data is a laborious process, and training data sets need to be updated on the appearance of new event log message types.
Thirdly, many existing template mining approaches have various design limitations. As pointed out in recent studies [26, 21], several algorithms rely on preprocessing event log messages, which requires previous domain knowledge of message formats. Also, as discussed in [26, 20], a number of algorithms incorrectly assume that the log messages described by the same template always contain the same number of words. When that is not the case (for example, two Suricata IDS alert messages from Figure 1 consist of 17 and 23 words, respectively), the algorithms fail to detect a common template for such messages.
Finally, existing studies have often used data sets which are either not recent or not related to cyber security. For example, data sets from the widely used Loghub repository [27] are not cyber security-centric, with several of them more than 15 years old.
@@
  • Whereas several existing works have utilized public LLMs [22, 20, 23], security event logs usually contain sensitive data that can not be shared with external service providers. For this reason, the current paper studies the performance of LLM-TD with several small local LLMs.
@@
  • The paper analyzes performance metrics that have been previously suggested for template detection algorithms [24, 25, 18], and proposes several heuristic principles for addressing the shortcomings of existing metrics.
  • Finally, the paper highlights the advantages of LLM-based template detection over traditional algorithms, and discusses the importance of qualitative assessment of all identified templates.
The remainder of this paper is organized as follows – Section II discusses related work, Section III describes the LLM-TD method, Section IV presents its evaluation, Section V discusses the evaluation results, and Section VI concludes the paper.
II Related Work
@@
SLCT [13] begins the template mining process by detecting frequent words from event log messages, continuing with creating cluster candidates based on the combinations of frequent words. Each candidate is represented by a template, and candidates that occur for at least s messages (s is a user-given threshold) are selected as clusters, with their templates being reported to the end user.
@@
LogCluster [12] is a further development of SLCT that employs an advanced cluster candidate generation procedure. That procedure allows us to address the shortcomings of many other algorithms discussed in Section I – templates detected by LogCluster can represent similar messages that contain different numbers of words. LFA [14] begins with identifying frequent words like SLCT and LogCluster and builds templates around frequent words that have similar occurrence times in event log messages.
@@
LenMa [11] relies on the following observation – the constant words of the template have the same length (number of characters) in all messages matching that template, whereas words matching the wildcards do not have that property. Following that observation, LenMa converts event log messages into word length vectors, assigning messages with the same number of words into the same cluster based on the cosine similarity of their word length vectors. Each cluster has a template representative which is reported to the end user.
@@
LKE [10] employs expert rules for removing variable parts from event log messages, and messages are then clustered based on their edit distance, deriving a template for each cluster. AEL [9] uses heuristic rules for identifying variable words in event log messages, dividing messages into bins based on the number of constant and variable words, and creating one or more templates for each bin. LogMine [15] first clusters event log messages by measuring their similarity in terms of common words and then generates a template for each cluster. These steps can be applied iteratively to produce more general templates from more specific ones.
@@
IPLoM [7] is a divisive clustering algorithm which first splits the event log into clusters by the number of words in log messages. For each cluster, a word position is then identified with the smallest number of unique words, splitting the cluster by these words. Resulting clusters are split further by associations between word pairs. After that step, a template is derived for each cluster.
@@
LogSig [6] converts each log message into word pairs and finds an optimal partition of messages into clusters, trying to maximize the number of common word pairs for messages in the same cluster. Spell [8] divides the event log messages based on the longest common subsequences (LCSs) of words, creating a new LCS object for a message if it is dissimilar to existing objects or merging it with the most similar object. Each LCS object contains information about the locations of variable words in log messages corresponding to that object. After all messages have been processed, templates that are based on LCS objects are reported to the end user.
@@
Logram [18] uses n-gram analysis for template detection, where each n-gram is a sequence of n words extracted from a log message. Logram begins with creating a dictionary of 2-grams and 3-grams, which holds their occurrence times. For template detection, Logram relies on the following observation – n-grams of constant words are frequent, while n-grams containing variable words are likely infrequent. To detect variable words, Logram first identifies infrequent 3-grams in event log messages and then analyzes infrequent 2-grams generated from infrequent 3-grams of the log message.
@@
Drain [5] employs a parse tree with a user-specified depth where templates are created and updated in leaf nodes. The parse tree arranges templates by the number of words in log messages and words at the beginning of messages. The similarity score between the message and the template is calculated to select the most suitable template for a log message, which has to exceed a user-given similarity threshold.
@@
SHISO [17] uses a tree structure for clustering log messages on the fly, and n-gram analysis for adjusting templates during log message processing. MoLFI [16] employs a genetic algorithm for generating templates from event log messages.
@@
Unlike previously described traditional algorithms, LLM-based template mining approaches are generally supervised. LogPPT [21] harnesses the Adaptive Random Sampling algorithm for selecting training data points from preprocessed event log data. Training data is then labeled with ground truth templates by the human analyst and used to fine-tune (train) a pretrained RoBERTa language model that will be applied for template detection. The authors evaluated the model performance on 2,000-message labeled event logs from the Loghub repository.
@@
LLMParser [19] employs four pretrained open-source LLMs fine-tuned with labeled examples selected from preprocessed event log data. For example selection, Mean Shift algorithm is used to cluster the preprocessed event log, and examples are sampled from detected clusters. The performance of models was measured on event logs of 2,000 messages from the Loghub repository. The authors also evaluated the performance of in-context learning, where instead of fine-tuning, a prompt (instructions in natural language) is presented to an LLM on how to address the template mining task. According to the results in [19], fine-tuning yielded higher performance than in-context learning.
@@
DivLog [20] uses in-context learning for template detection, harnessing the public GPT-3 LLM through the OpenAI interface. The algorithm samples a subset of event log data with the Determinantal Point Process algorithm, and selected log messages are labeled by the human analyst. To detect a template for the log message, DivLog builds a prompt of similar message-template pairs selected from the labeled data so that LLM can learn from the provided examples to find the right template for the given message. The performance of DivLog was evaluated on 2,000-message event logs from Loghub.
@@
LILAC [22] uses in-context learning with the public ChatGPT LLM through the OpenAI interface, where the human analyst needs to label a subset of event log messages for prompt-building purposes. For selecting messages for labeling, a hierarchical clustering algorithm is applied for the event log, sampling messages from different clusters based on a quota distribution method. When LILAC needs to detect a template for the log message, it first checks its template cache for a suitable template in order to minimize queries to LLM. Such a design allows to apply LILAC for processing large event logs, and the authors evaluated the performance of LILAC on full-size event logs (i.e., not 2,000-message event log chunks as in several other works [21, 19, 20]) from Loghub.
@@
Similarly to LILAC, ULog [23] employs ChatGPT LLM through OpenAI interface. Unlike previously discussed LLM-based algorithms which are supervised [21, 19, 20, 22], ULog is an unsupervised method and does not rely on example log messages labeled by a human analyst with ground truth templates. Whereas other approaches select specific labeled log message examples for building prompts, so that a template could be detected for one log message according to labeled examples, ULog employs a different heuristic – if several similar log messages are submitted to LLM, LLM is able to find a common template for them without labeled examples. Ideally, log message groups submitted to LLM should match the same template, while being diverse enough so that LLM can distinguish variable parts from constant parts in log messages. For example, in the first message group in Figure 1, all variable parts of two log messages are different, which eases the template detection task for LLM.
@@
ULog employs a hierarchical log clustering algorithm to detect groups of similar log messages. First, like many traditional algorithms, ULog assumes that log messages matching the same template contain the same number of words, and the event log is thus split into buckets by the number of words in log messages. By the design of ULog, the further processing of each bucket is independent of other buckets, which allows the execution of these tasks in parallel. In each bucket, top-k frequent words are identified, creating clusters for log messages that share the same frequent words, and merging clusters with too few log messages based on top-(k-1) frequent words. To select message groups for LLM, ULog samples so-called log contrastive units (LCUs) from each cluster, so that log messages in LCUs would be diverse enough. If a template is successfully detected for an LCU, all messages matching this template are removed from buckets. The sampling of LCUs from buckets and querying a common template for them from LLM continues until all buckets are empty.
III Unsupervised LLM-based Template Detection
@@
In this section, we will describe an unsupervised LLM-based template detection method (henceforth called LLM-TD) that has been motivated by research gaps in existing literature. As discussed in Section II, unsupervised LLM-based template detection has received little attention, apart from a recent work proposing ULog [23]. However, ULog suffers from a design limitation discussed in Section I, assuming that log messages matching the same template have the same number of words (see Section II for more details).
@@
Secondly, several existing works [23, 22, 20] employ public LLMs through the OpenAI interface, which implies submitting event log data to an external service provider. However, security logs are likely to contain highly sensitive information (e.g., information about user accounts and authentication methods in an organization), and sharing it with external parties is regarded as a violation of security policies.
@@
Whereas local LLMs allow processing event logs without external service providers, such LLMs are known to be less performant than public LLMs. For this reason, their use has been primarily considered with fine-tuning [21, 19], which, according to a recent study, yields better performance than in-context learning [19]. However, we argue that fine-tuning requires a considerable amount of machine learning expertise, which is not always present in an organization.
@@
For the reasons above, this paper explores the use of local LLMs for security event log analysis, and the LLM-TD method employs an unsupervised in-context learning paradigm for template detection. Whereas other methods attempt to detect one template with each LLM query (i.e., templates are detected one-by-one), LLM-TD uses a different approach, which is able to identify several templates with one query. In order to achieve that, LLM-TD utilizes a prompt that formulates the task of a template detection for a small event log which contains messages matching several templates, not for one message or a group of similar messages which match one template only.
@@
When compared to ULog, which is another unsupervised LLM-based algorithm (see Section II), both LLM-TD and ULog process a batch of several log messages with one LLM query because it can ease the unsupervised template detection task for LLMs [23]. However, ULog harnesses a dedicated multi-stage selection method for ensuring the similarity of messages within a batch, so that a single template can be detected for the batch. In contrast, as LLM-TD does not aim to detect one template which must cover the entire batch, it can process less similar messages with one LLM query, and thus does not need a complex method for creating message batches.
@@
As pointed out in [23], unsupervised detection of templates is complicated if too diverse log messages are submitted to LLM. Following that observation, LLM-TD divides the event log data into different parts by applications (e.g., sshd, Suricata) that have produced the log messages, where each part is processed independently. Similarly to ULog, dividing the event log data allows for parallel processing but does not restrict templates to match the messages with the same number of words (unlike ULog). Having each part containing messages from the same application leads to much less diversity among log messages and fewer templates that need detection.
@@
Also, in the case of Linux syslog-based security logs (the primary focus of this paper), the application name can be retrieved from the syslog tag field which is defined by the BSD syslog protocol [4]. Since the purpose of LLM-TD is the analysis of unstructured log message data, LLM-TD implementation will not consider timestamp and hostname fields of syslog messages, targeting the syslog tag field which is followed by a free-form textual message string (Figure 1 depicts such syslog message examples).
@@
Figure 2: Using an LLM for detecting templates from an event log.
@@
IV Evaluation
IV-A Experiment Environment and Data Sets
For conducting the experiments with LLM-TD on Linux security syslog files, we created an implementation of LLM-TD in Perl, with the QueryLLMforTemplates procedure from Algorithm 1 being implemented in Python using the LangChain framework. For local LLMs, we selected several local LLMs – OpenChat [28], Mistral [29], and Wizardlm2 [30]. In the case of some benchmarks [28], OpenChat has been reported to offer comparable performance to ChatGPT, which has been employed by several previous works for template detection [22, 23]. Also, Mistral and Wizardlm2 are state-of-the-art LLMs from Mistral AI and Microsoft AI, respectively, which we found to work well for template detection tasks. For all three models, we employed their 7B implementations through the Ollama framework, which was installed on a Rocky Linux 9 workstation with NVIDIA GeForce RTX 2080 Ti video card, Intel Core i9-10900X CPU, and 128GB of memory. Therefore, the evaluation allowed us to establish how well small LLMs with parameter size 7B would be suited for mining templates from security event logs using the computing power of only one workstation.
In order to find how well LLM-TD with small LLMs is performing in comparison to traditional algorithms, we selected Drain [5] as a baseline because it was identified as the most performant traditional algorithm in the study by Zhu et al. [24]. Also, Drain has been used as a baseline in several recent studies on LLM-based template detection [21, 20, 22, 23, 19]. We were not able to include ULog (the only existing unsupervised LLM-based algorithm) in the comparison since its implementation is not publicly available.
TABLE I: Event Log Data Sets

### v3

@@ Abstract
journal: International Journal of Information Security
1 Introduction
Event log analysis is an important cyber security monitoring technique. For example, modern Security Operations Center (SOC) platforms like Splunk splunk and ElasticStack elasticstack rely on collecting and analyzing event logs from a large number of sources, whereas dedicated event correlation tools like SEC sec have been designed for real-time monitoring of event logs. Although during recent years several structured event log formats have been proposed, a significant part of event log messages remain unstructured or semi-structured. As an example, the widely used BSD syslog protocol bsd-syslog defines the priority, timestamp, hostname, and syslog tag (program name with its process ID) fields, whereas the message text, which is the most informative part of the syslog message, remains a free-form string. Figure 1 depicts some example syslog messages without the priority, timestamp, and hostname fields (note that all syslog messages have been wrapped in Figure 1 due to their length, and they appear in one line in real-life event logs).
Figure 1: Example syslog messages and their templates
@@ 1 Introduction
During the last two decades, a number of template detection algorithms have been proposed drain ; logsig ; iplom ; spell ; ael ; lenma ; logcluster ; slct ; lfa ; lke ; logmine ; shiso ; molfi ; logram ; logppt ; divlog ; lilac ; llmparser ; ulog , and several comparative studies have been published for these algorithms khan ; zhu ; gasimov . Despite a significant amount of research in the domain of template mining, some research gaps have remained. Firstly, template detection approaches that have not employed LLMs (we call them traditional algorithms in the remainder of this paper) are unsupervised data mining-based approaches which can be tuned through hyperparameters. However, as pointed out in logppt , selecting optimal values for hyperparameters is often not an easy task. Secondly, LLM-based template detection approaches are generally supervised and require labeled training data llmparser ; divlog ; lilac ; logppt . As discussed in ulog , the creation of such training data is a laborious process, and training data sets need to be updated on the appearance of new event log message types.
Thirdly, many existing template mining approaches have various design limitations. As pointed out in recent studies gasimov ; logppt , several algorithms rely on preprocessing event log messages, which requires previous domain knowledge of message formats. Also, as discussed in gasimov ; divlog , a number of algorithms incorrectly assume that the log messages described by the same template always contain the same number of words. When that is not the case (for example, two Suricata IDS alert messages from Figure 1 consist of 17 and 23 words, respectively), the algorithms fail to detect a common template for such messages.
Finally, existing studies have often used data sets which are either not recent or not related to cyber security. For example, data sets from the widely used Loghub repository loghub are not cyber security-centric, with several of them more than 15 years old.
@@
  • Whereas several existing works have utilized public LLMs lilac ; divlog ; ulog , security event logs usually contain sensitive data that can not be shared with external service providers. For this reason, the current paper studies the performance of LLM-TD with several small local LLMs.
@@
  • The paper analyzes performance metrics that have been previously suggested for template detection algorithms zhu ; khan ; logram , and proposes several heuristic principles for addressing the shortcomings of existing metrics.
  • Finally, the paper highlights the advantages of LLM-based template detection over traditional algorithms, and discusses the importance of qualitative assessment of all identified templates. In particular, LLMs have the ability to infer correct templates from insufficient event log data and identify character patterns inside log message words. Furthermore, they are able to discover previously unknown knowledge from event logs, making the qualitative analysis of both correct and incorrect templates a relevant task.
The remainder of this paper is organized as follows – Section 2 discusses related work, Section 3 describes the LLM-TD method, Section 4 presents its evaluation, Section 5 discusses the evaluation results, and Section 6 concludes the paper.
2 Related Work
@@
SLCT slct begins the template mining process by detecting frequent words from event log messages, continuing with creating cluster candidates based on the combinations of frequent words. Each candidate is represented by a template, and candidates that occur for at least s messages (s is a user-given threshold) are selected as clusters, with their templates being reported to the end user.
@@
LogCluster logcluster is a further development of SLCT that employs an advanced cluster candidate generation procedure. That procedure allows us to address the shortcomings of many other algorithms discussed in Section 1 – templates detected by LogCluster can represent similar messages that contain different numbers of words. LFA lfa begins with identifying frequent words like SLCT and LogCluster and builds templates around frequent words that have similar occurrence times in event log messages.
@@
LenMa lenma relies on the following observation – the constant words of the template have the same length (number of characters) in all messages matching that template, whereas words matching the wildcards do not have that property. Following that observation, LenMa converts event log messages into word length vectors, assigning messages with the same number of words into the same cluster based on the cosine similarity of their word length vectors. Each cluster has a template representative which is reported to the end user.
@@
LKE lke employs expert rules for removing variable parts from event log messages, and messages are then clustered based on their edit distance, deriving a template for each cluster. AEL ael uses heuristic rules for identifying variable words in event log messages, dividing messages into bins based on the number of constant and variable words, and creating one or more templates for each bin. LogMine logmine first clusters event log messages by measuring their similarity in terms of common words and then generates a template for each cluster. These steps can be applied iteratively to produce more general templates from more specific ones.
@@
IPLoM iplom is a divisive clustering algorithm which first splits the event log into clusters by the number of words in log messages. For each cluster, a word position is then identified with the smallest number of unique words, splitting the cluster by these words. Resulting clusters are split further by associations between word pairs. After that step, a template is derived for each cluster.
@@
LogSig logsig converts each log message into word pairs and finds an optimal partition of messages into clusters, trying to maximize the number of common word pairs for messages in the same cluster. Spell spell divides the event log messages based on the longest common subsequences (LCSs) of words, creating a new LCS object for a message if it is dissimilar to existing objects or merging it with the most similar object. Each LCS object contains information about the locations of variable words in log messages corresponding to that object. After all messages have been processed, templates that are based on LCS objects are reported to the end user.
@@
Logram logram uses n-gram analysis for template detection, where each n-gram is a sequence of n words extracted from a log message. Logram begins with creating a dictionary of 2-grams and 3-grams, which holds their occurrence times. For template detection, Logram relies on the following observation – n-grams of constant words are frequent, while n-grams containing variable words are likely infrequent. To detect variable words, Logram first identifies infrequent 3-grams in event log messages and then analyzes infrequent 2-grams generated from infrequent 3-grams of the log message.
@@
Drain drain employs a tree structure called parse tree with a user-specified depth where templates are created and updated in leaf nodes. For all log messages with the same number of words, a separate child node is created under the root node of the parse tree. The nodes at the following levels contain message words and wildcards, allowing to search for a leaf node with a list of relevant templates for each log message. The similarity score between the message and the template is calculated to select the most suitable template for a log message, which has to exceed a user-given similarity threshold, and the selected template is updated (if no suitable template is found, a new template created).
@@
SHISO shiso uses a tree structure for clustering log messages on the fly, and n-gram analysis for adjusting templates during log message processing. MoLFI molfi employs a genetic algorithm for generating templates from event log messages.
@@
Unlike previously described traditional algorithms, LLM-based template mining approaches are generally supervised. LogPPT logppt harnesses the Adaptive Random Sampling algorithm for selecting training data points from preprocessed event log data. Training data is then labeled with ground truth templates by the human analyst and used to fine-tune (train) a pretrained RoBERTa language model that will be applied for template detection. The authors evaluated the model performance on 2,000-message labeled event logs from the Loghub repository.
@@
LLMParser llmparser employs four pretrained open-source LLMs fine-tuned with labeled examples selected from preprocessed event log data. For example selection, Mean Shift algorithm is used to cluster the preprocessed event log, and examples are sampled from detected clusters. The performance of models was measured on event logs of 2,000 messages from the Loghub repository. The authors also evaluated the performance of in-context learning, where instead of fine-tuning, a prompt (instructions in natural language) is presented to an LLM on how to address the template mining task (please see Figure 2 in Section 3 for an example prompt). According to the results in llmparser , fine-tuning yielded higher performance than in-context learning. As discussed in llmparser , used LLMs accepted prompts of limited size, and insufficient log parsing instructions with a very few examples had a negative impact on performance. In contrast, fine-tuning enabled the model to learn from a larger number of more diverse examples, yielding a better performance.
@@
DivLog divlog uses in-context learning for template detection, harnessing the public GPT-3 LLM through the OpenAI interface. The algorithm samples a subset of event log data with the Determinantal Point Process algorithm, and selected log messages are labeled by the human analyst. To detect a template for the log message, DivLog builds a prompt of similar message-template pairs selected from the labeled data so that LLM can learn from the provided examples to find the right template for the given message. The performance of DivLog was evaluated on 2,000-message event logs from Loghub.
@@
LILAC lilac uses in-context learning with the public ChatGPT LLM through the OpenAI interface, where the human analyst needs to label a subset of event log messages for prompt-building purposes. For selecting messages for labeling, a hierarchical clustering algorithm is applied for the event log, sampling messages from different clusters based on a quota distribution method. When LILAC needs to detect a template for the log message, it first checks its template cache for a suitable template in order to minimize queries to LLM. Such a design allows to apply LILAC for processing large event logs, and the authors evaluated the performance of LILAC on full-size event logs (i.e., not 2,000-message event log chunks as in several other works logppt ; llmparser ; divlog ) from Loghub.
@@
Similarly to LILAC, Lunar ulog employs ChatGPT LLM through OpenAI interface. Unlike previously discussed LLM-based algorithms which are supervised logppt ; llmparser ; divlog ; lilac , Lunar is an unsupervised method and does not rely on example log messages labeled by a human analyst with ground truth templates. Whereas other approaches select specific labeled log message examples for building prompts, so that a template could be detected for one log message according to labeled examples, Lunar employs a different heuristic – if several similar log messages are submitted to LLM, LLM is able to find a common template for them without labeled examples. Ideally, log message groups submitted to LLM should match the same template, while being diverse enough so that LLM can distinguish variable parts from constant parts in log messages. For example, in the first message group in Figure 1, all variable parts of two log messages are different, which eases the template detection task for LLM.
@@
Lunar employs a hierarchical log clustering algorithm to detect groups of similar log messages. First, like many traditional algorithms, Lunar assumes that log messages matching the same template contain the same number of words, and the event log is thus split into buckets by the number of words in log messages. By the design of Lunar, the further processing of each bucket is independent of other buckets, which allows the execution of these tasks in parallel. In each bucket, top-k frequent words are identified, creating clusters for log messages that share the same frequent words, and merging clusters with too few log messages based on top-(k-1) frequent words. To select message groups for LLM, Lunar samples so-called log contrastive units (LCUs) from each cluster, so that log messages in LCUs would be diverse enough. If a template is successfully detected for an LCU, all messages matching this template are removed from buckets. The sampling of LCUs from buckets and querying a common template for them from LLM continues until all buckets are empty.
3 Unsupervised LLM-based Template Detection
@@
In this section, we will describe an unsupervised LLM-based template detection method (henceforth called LLM-TD) that has been motivated by research gaps in existing literature. As discussed in Section 2, unsupervised LLM-based template detection has received little attention, apart from a recent work proposing Lunar ulog . However, Lunar suffers from a design limitation discussed in Section 1, assuming that log messages matching the same template have the same number of words (see Section 2 for more details). Because unsupervised template detection does not need a larger pool of log message examples labeled with ground truth templates, it allows to eliminate the labeling workload of human analysts, and LLM-TD is therefore employing this template detection paradigm.
@@
Secondly, several existing works ulog ; lilac ; divlog employ public LLMs through the OpenAI interface, which implies submitting event log data to an external service provider. However, security logs are likely to contain highly sensitive information (e.g., information about user accounts and authentication methods in an organization), and sharing it with external parties is regarded as a violation of security policies.
@@
Whereas local LLMs allow processing event logs without external service providers, such LLMs are known to be less performant than public LLMs. For this reason, their use has been primarily considered with fine-tuning logppt ; llmparser , which, according to a recent study, yields better performance than in-context learning llmparser . However, we argue that fine-tuning requires a considerable amount of machine learning expertise, which is not always present in an organization.
@@
For the reasons above, this paper explores the use of local LLMs for security event log analysis, and the LLM-TD method employs an unsupervised in-context learning paradigm for template detection. Whereas other methods attempt to detect one template with each LLM query (i.e., templates are detected one-by-one), LLM-TD uses a different approach, which is able to identify several templates with one query. In order to achieve that, LLM-TD utilizes a prompt that formulates the task of a template detection for a small event log which contains messages matching several templates, not for one message or a group of similar messages which match one template only (Figure 2 depicts the prompt used by LLM-TD).
@@
When compared to Lunar, which is another unsupervised LLM-based algorithm (see Section 2), both LLM-TD and Lunar process a batch of several log messages with one LLM query because it can ease the unsupervised template detection task for LLMs ulog . However, Lunar harnesses a dedicated multi-stage selection method for ensuring the similarity of messages within a batch, so that a single template can be detected for the batch. In contrast, as LLM-TD does not aim to detect one template only which must cover the entire batch, it can process less similar messages with one LLM query, and thus does not need a complex method for creating message batches. Note that LLM-TD does not specify any constraints to the number of detected templates in LLM query, and LLM is free to report back any number of templates (e.g., one, two, five, etc.) as it sees most appropriate for the submitted batch.
@@
As pointed out in ulog , unsupervised detection of templates is complicated if too diverse log messages are submitted to LLM. Following that observation, LLM-TD divides the event log data into different parts by applications (e.g., sshd, Suricata) that have produced the log messages, where each part is processed independently. Similarly to Lunar, dividing the event log data allows for parallel processing but does not restrict templates to match the messages with the same number of words (unlike Lunar). Having each part containing messages from the same application leads to much less diversity among log messages and fewer templates that need detection.
@@
Also, in the case of Linux syslog-based security logs (the primary focus of this paper), the application name can be retrieved from the syslog tag field which is defined by the BSD syslog protocol bsd-syslog . Since the purpose of LLM-TD is the analysis of unstructured log message data, LLM-TD implementation will not consider timestamp and hostname fields of syslog messages, targeting the syslog tag field which is followed by a free-form textual message string (Figure 1 depicts such syslog message examples). Note that the location of timestamp, hostname, syslog tag (application name), and message string fields are well documented in the syslog protocol standard bsd-syslog , and therefore the identification of these fields does not assume detailed domain knowledge about the individual message formats of logging applications.
@@
Figure 2: Interaction between LLM-TD and the underlying LLM (QueryLLMforTemplates and Merge procedures from Algorithm 1)
@@
4 Evaluation
4.1 Experiment Environment and Data Sets
For conducting the experiments with LLM-TD on Linux security syslog files, we created an implementation of LLM-TD in Perl, with the QueryLLMforTemplates procedure from Algorithm 1 being implemented in Python using the LangChain framework. For local LLMs, we selected three local LLMs – OpenChat openchat , Mistral mistral , and Wizardlm2 wizardlm2 . In the case of some benchmarks openchat , OpenChat has been reported to offer comparable performance to ChatGPT, which has been employed by several previous works for template detection lilac ; ulog . Also, Mistral and Wizardlm2 are state-of-the-art LLMs from Mistral AI and Microsoft AI, respectively, which we found to work well for template detection tasks. We also considered a number of other local LLMs for general purposes (Llama2, Llama3, Mixtral, Phi, and Phi3) and more specific goals such as code analysis and mathematical reasoning (Codellama, Codeup, and Wizard-math), but they featured lower runtime and template detection performance than the selected three LLMs.
For all three models, we employed their 7B implementations through the Ollama framework, which was installed on a Rocky Linux 9 workstation with NVIDIA GeForce RTX 2080 Ti video card, Intel Core i9-10900X CPU, and 128GB of memory. Therefore, the evaluation allowed us to establish how well small LLMs with parameter size 7B would be suited for mining templates from security event logs using the computing power of only one workstation.
In order to find how well LLM-TD with small LLMs is performing in comparison to traditional algorithms, we selected Drain drain as a baseline because it was identified as the most performant traditional algorithm in the study by Zhu et al. zhu . Also, Drain has been used as a baseline in several recent studies on LLM-based template detection logppt ; divlog ; lilac ; ulog ; llmparser . We were not able to include Lunar (the only existing unsupervised LLM-based algorithm) in the comparison since its implementation is not publicly available.
Table 1: Event log data sets